# The Effect of Personalization in FedProx: A Fine-grained Analysis on Statistical Accuracy and Communication Efficiency

### v3

@@ 7 Conclusion
In this paper, we analyzed of the effect of personalization in FedProx. With a theoretically-informed adaptive choice of personalization degree, FedProx achieves a minimax-optimal statistical accuracy under different statistical heterogeneity degrees and consistently outperforms pure local training. As a minor contribution, we provably show that personalization reduces communication costs without increasing the computation overhead. The results are validated through extensive experiments.
@@ Appendix A Proof of statistical converngence
where $\tilde{\Delta}^{(g)}_{\text{stat}}$ is the difference between the true global model and the optimal solution of the global model in (1) and $\tilde{\Delta}^{(i)}_{\text{stat}}$ is the difference between the true local model and the optimal solution of the local model in (1) for any $i\in[m]$. The statistical error bound is established as $\mathbb{E}\|\tilde{\Delta}^{(g)}_{\text{stat}}\|^{2}$ and $\mathbb{E}\|\tilde{\Delta}^{(i)}_{\text{stat}}\|^{2}$, where the expectation is taken over all the data.
And we also denote
@@
A.1 Discussion on Assumption 1
When assuming the statistical heterogeneity of each client is different ($R$ is related to the client index $i$), it would provide a more delicate description of statistical heterogeneity, and it also requires different personalization degree per client. Therefore, it is better to consider the objective (18), where the $i$-th client will be shrunk to the global with strength $\lambda_{i}$. Therefore, solving (18) will consider different personalization degree for different clients. Solving the objective (18) would be open for future work.
@@ A.1 Discussion on Assumption 1
Under Assumptions 2 and 3 as we assume, Assumption 1 about statistical heterogeneity would be equivalent to many other similar assumptions for statistical heterogeneity, like $B$-dissimilarity (Li et al., 2020a), parameter difference (Chen et al., 2023b) and gradient diversity (T Dinh et al., 2020; Deng et al., 2020).
A.2 Lemmas and the Proof
To facilitate the proof of the theorem, we first present two additional lemmas and provide their proof.
@@ A.2 Lemmas and the Proof
where in the last step we used the Cauchy inequality.
In addition, notice that ${\tilde{\bm{w}}^{(i)}}$ is the optimal solution for the local model ${{\bm{w}}^{(i)}}$ in the FedProx problem (1), thus we have
@@
as claimed.
@@
Under Assumption 1, 2, 3, 4, for $\tilde{\bm{w}}_{\text{GT}}$, the solution of the GlobalTrain problem in (2), we have
@@
$\leq\frac{\Big{\|}\sum_{i\in[m]}p_{i}\nabla L_{i}({{\bm{w}}_{\star}^{(i)}},S_{i})\Big{\|}+\sum_{i\in[m]}p_{i}\Big{\|}\nabla L_{i}({{\bm{w}}_{\star}^{(g)}},S_{i})-\nabla L_{i}({{\bm{w}}_{\star}^{(i)}},S_{i})\Big{\|}}{\mu/2}$ (21)
@@
as claimed.
A.3 Proof of Lemma 1 and Theorem 1
A.3.1 Proof of Lemma 1
Lemma 1 is a direct consequence of the proof of Theorem 1. To see this, from (A.3.2) and the fact that $(\mu+\lambda)^{2}\geq\mu(\mu+\lambda)$, we have
$\|{\tilde{\bm{w}}^{(g)}}-\tilde{\bm{w}}_{\text{GT}}\|^{2}$ $\leq\frac{4}{\mu}\left(\frac{1}{\mu+\lambda}+\frac{L}{(\mu+\lambda)^{2}}\right)\sum_{i}p_{i}\|\nabla L_{i}({\tilde{\bm{w}}^{(g)}},S_{i})\|^{2}$
$\leq\frac{4(1+L/\mu)}{\mu(\mu+\lambda)}\sum_{i}p_{i}\|\nabla L_{i}({\tilde{\bm{w}}^{(g)}},S_{i})\|^{2},$
thus (4) holds. In addition, with the fact that $\|\tilde{\bm{\delta}}^{(i)}\|_{2}\geq 0$, from (28) we find
00 0 $\geq-\sum_{i}\sum_{j}\frac{p_{i}}{n_{i}}\left\langle\nabla\ell({{\bm{w}}_{\star}^{(i)}},z_{ij}),\tilde{\Delta}^{(i)}_{\text{stat}}\right\rangle+\frac{\mu}{2}\sum_{i}\sum_{j}\frac{p_{i}}{n_{i}}\left\|\tilde{\Delta}^{(i)}_{\text{stat}}\right\|^{2}+\frac{\lambda}{2}\sum_{i}p_{i}\left\|\tilde{\bm{\delta}}^{(i)}\right\|^{2}-\frac{\lambda}{2}R^{2}$
$\geq-\sum_{i}\sum_{j}\frac{p_{i}}{n_{i}}\left\langle\nabla\ell({{\bm{w}}_{\star}^{(i)}},z_{ij}),\tilde{\Delta}^{(i)}_{\text{stat}}\right\rangle+\frac{\mu}{2}\sum_{i}\sum_{j}\frac{p_{i}}{n_{i}}\left\|\tilde{\Delta}^{(i)}_{\text{stat}}\right\|^{2}-\frac{\lambda}{2}R^{2}.$
which implies the result in (5).
A.3.2 Proof of Global Statistical Error Bound in Theorem 1
@@
Recall that ${\tilde{\bm{w}}^{(g)}}$ and $\{{\tilde{\bm{w}}^{(i)}}\}_{i\in[m]}$ are the minimizers of the objective (1). According to the first-order condition, we have
@@
For the first term on the RHS in (25), apply the $\mu$-strongly convexity of the loss function $\ell$ (cf. Assumption 3), we then have
@@
Recall the definition in (17) and the Assumption 1 about statistical heterogeneity
@@
The equation (33) would be regarded as a quadratic function of $\sum_{i}p_{i}\mathbb{E}\left\|\tilde{\Delta}^{(i)}_{\text{stat}}\right\|^{2}$. If we set the $\lambda$ as
$\lambda\leq\frac{\rho^{2}}{\mu R^{2}}\sum_{i\in[m]}\frac{p_{i}}{n_{i}},$ (34)
then solving the inequality in (33) w.r.t $\sum_{i}p_{i}\mathbb{E}\left\|\tilde{\Delta}^{(i)}_{\text{stat}}\right\|^{2}$ yields
$\begin{split}\mathbb{E}\left\|\tilde{\Delta}^{(g)}_{\text{stat}}\right\|_{2}&\leq\left(\sum_{i}p_{i}\mathbb{E}\left\|\tilde{\Delta}^{(i)}_{\text{stat}}\right\|^{2}\right)^{\frac{1}{2}}\leq\frac{1}{\mu}\left(\rho\left(\sum_{i\in[m]}\frac{p_{i}}{n_{i}}\right)^{\frac{1}{2}}+\left(\rho^{2}\sum_{i\in[m]}\frac{3p_{i}}{n_{i}}\right)^{\frac{1}{2}}\right)\\ &=\frac{(1+\sqrt{3})\rho}{\mu}\left(\sum_{i\in[m]}\frac{p_{i}}{n_{i}}\right)^{\frac{1}{2}}.\\ \end{split}$ (35)
@@
$\sum_{i}p_{i}\nabla L_{i}(\tilde{\bm{w}}_{\text{GT}},S_{i})=0,$ (42)
@@
Take expectation w.r.t $S_{i}$ and use Assumption 1 and 4, then we have
@@
If we assume
we have
Furthermore, we can get a simplified condition for $\lambda$ as
$\lambda\geq\frac{8\kappa}{\mu}\left\{8L^{2}\vee\left(\left(2L^{2}R^{2}+\frac{\rho^{2}}{n}\right)\Big{/}\left(\frac{4\rho^{2}}{\mu^{2}N}+\frac{4L^{2}R^{2}}{\mu^{2}}\right)\right)\right\}-\mu$ (54)
Combine two-case arguments, we have
$\Rightarrow\mathbb{E}\|{\tilde{\bm{w}}^{(g)}}-{{\bm{w}}_{\star}^{(g)}}\|^{2}\leq\begin{cases}\left(\frac{(1+\sqrt{3})\rho}{\mu}\right)^{2}\sum_{i\in[m]}\frac{p_{i}}{n_{i}}&R>\sqrt{\sum_{i}\frac{p_{i}}{n_{i}}}\\ 32\left(\frac{\rho^{2}}{\mu^{2}}\frac{1}{N}+\frac{L^{2}}{\mu^{2}}R^{2}\right)&R\leq\sqrt{\sum_{i}\frac{p_{i}}{n_{i}}}.\end{cases}$ (55)
Thus it yields a one-line rate
$\mathbb{E}\|{\tilde{\bm{w}}^{(g)}}-{{\bm{w}}_{\star}^{(g)}}\|^{2}\leq C_{1}\frac{1}{N}+C_{2}(\sum_{i\in[m]}\frac{p_{i}}{n_{i}}\wedge R^{2}),$ (56)
where
Moreover, if we assume data balanced, i.e $n_{i}=n,\forall i\in[m]$ and $p_{i}=\frac{1}{m}$, we obtain
$\mathbb{E}\|{\tilde{\bm{w}}^{(g)}}-{{\bm{w}}_{\star}^{(g)}}\|^{2}\leq C_{1}\frac{1}{N}+C_{2}\left(\frac{1}{n}\wedge R^{2}\right),$ (58)
as claimed.
A.3.3 Proof of Local Statistical Error Bound in Theorem 1
@@
For the first term on the RHS, applying the $\mu$-strongly convex of the loss function $\ell$ (cf. Assumption 3) yields
@@
Pluggging (61) back into (60), it yields
@@
To utilize the statistical global convergence rate shown in Appendix A.3.2, we need to discuss $R$ in the same way. Recall the upper bound of $\tilde{\Delta}^{(g)}_{\text{stat}}$ in (58):
$\mathbb{E}\|\tilde{\Delta}^{(g)}_{\text{stat}}\|^{2}\leq\begin{cases}\frac{C_{2}}{n}&R>\frac{1}{\sqrt{n}}\\ \frac{C_{1}}{N}+C_{2}R^{2}&R\leq\frac{1}{\sqrt{n}}.\end{cases}$ (67)
For simplicity, in the following arguments, we denote the upper bound of $\mathbb{E}\|\tilde{\Delta}^{(g)}_{\text{stat}}\|_{2}$ as $U_{0}$.
@@
Putting (67) into (69) and reorganizing, we obtain
@@
In addition, recall that in this case we set $\lambda^{\star}\leq\frac{\rho^{2}}{\mu R^{2}}\sum_{i\in[m]}\frac{p_{i}}{n_{i}}$ in the statistical convergence analysis of the global model. If we assume data balanced, i.e. $p_{i}=\frac{1}{m}$ and $n_{i}=n,\forall i\in[m]$, it yields $\lambda^{\star}\leq\frac{\rho^{2}}{\mu nR^{2}}$ and
$\begin{split}\mathbb{E}\|\tilde{\Delta}^{(i)}_{\text{stat}}\|_{2}&\leq\frac{1}{\mu}\left[\frac{\rho}{\sqrt{n}}+\left(\frac{\rho^{2}}{n}+\frac{2\rho^{2}}{nR^{2}}\left(U_{0}+R\right)^{2}\right)^{\frac{1}{2}}\right].\\ &=\frac{1}{\sqrt{n}}\frac{1}{\mu}\left(\rho+\left(\rho^{2}+\frac{2\rho^{2}}{R^{2}}(U_{0}+R)^{2}\right)^{\frac{1}{2}}\right)\\ &\leq\frac{1}{\sqrt{n}}\frac{\rho+(\rho^{2}+4\rho^{2}(C_{2}+1))^{\frac{1}{2}}}{\mu}\\ &=\frac{1}{\sqrt{n}}\frac{\rho(1+(1+4(C_{2}+1))^{\frac{1}{2}})}{\mu}.\end{split}$ (71)
@@
$\overset{(b)}{\leq}\frac{12}{(\mu+\lambda)^{2}}\left[4L^{2}\mathbb{E}[\|{\tilde{\bm{w}}^{(g)}}-{{\bm{w}}_{\star}^{(g)}}\|^{2}]+4L^{2}R^{2}+2\frac{\rho^{2}}{n}\right]+3\mathbb{E}[\|{\tilde{\bm{w}}^{(g)}}-{{\bm{w}}_{\star}^{(g)}}\|^{2}]+3R^{2}$
$\overset{(c)}{\leq}\frac{12}{(\mu+\lambda)^{2}}\left[128L^{2}\left(\frac{\rho^{2}}{\mu^{2}}\frac{1}{N}+\frac{L^{2}}{\mu^{2}}R^{2}\right)+4L^{2}R^{2}+2\frac{\rho^{2}}{n}\right]+96\left(\frac{\rho^{2}}{\mu^{2}}\frac{1}{N}+\frac{L^{2}}{\mu^{2}}R^{2}\right)+3R^{2}$
@@
$=\frac{12}{(\mu+\lambda)^{2}}\left[\frac{128L^{2}\rho^{2}}{\mu^{2}}\frac{1}{N}+(128L^{4}/\mu^{2}+4L^{2})R^{2}+2\frac{\rho^{2}}{n}\right]+\frac{96\rho^{2}}{\mu^{2}}\frac{1}{N}+(96\frac{L^{2}}{\mu^{2}}+3)R^{2}$ (72)
where step (a) comes from Assumption 1 and Lemma 4, step (b) comes from result (A.3.2) and step (c) is a direct result of applying the error bound of $\mathbb{E}\|\tilde{\Delta}^{(g)}_{\text{stat}}\|^{2}$.

### v4

@@ 7 Conclusion
In this paper, we provide a precise theoretical characterization of the statistical and optimization convergence of a widely used personalized federated learning problem. Our analysis reveals that when collaborative learning is beneficial, increasing personalization reduces communication complexity but comes at the cost of statistical accuracy due to limited information sharing across clients. We then validate our theoretical findings across convex and non-convex settings, multiple datasets, and different model architectures.
Acknowledgment
This work was supported in part by the U.S. National Science Foundation (NSF) grant CCF-2007823 and National Institutes of Health (NIH) grants R01AI170249 and R01GM152812.
Impact Statement
This paper aims to contribute to the advancement of the field of Machine Learning. While our work has the potential for various societal implications, we believe none require specific emphasis in this context.
@@ Appendix A Proof of statistical converngence
where ${{\bm{w}}_{\star}^{(g)}}:=\sum_{i\in[m]}p_{i}{{\bm{w}}_{\star}^{(i)}}$, $\tilde{\Delta}^{(g)}_{\text{stat}}$ is the difference between the true global model and the optimal solution of the global model in Problem (2) and $\tilde{\Delta}^{(i)}_{\text{stat}}$ is the difference between the true local model and the optimal solution of the local model in (2) for any $i\in[m]$. The statistical error bound is established as $\mathbb{E}\|\tilde{\Delta}^{(g)}_{\text{stat}}\|^{2}$ and $\mathbb{E}\|\tilde{\Delta}^{(i)}_{\text{stat}}\|^{2}$, where the expectation is taken over all the data. Notably, we should stress that the definition of ${{\bm{w}}_{\star}^{(g)}}$ is just to serve as a bridge to establish the convergence rate of local models, and it is not our focus in the theoretical analysis.
We also denote
@@
A.1 Discussion on the Parameter Space
When assuming the statistical heterogeneity of each client is different ($R$ is related to the client index $i$), it would provide a more delicate description of statistical heterogeneity, and it also requires different personalization degree per client. Therefore, let’s consider the following problem with different heterogeneity in personalized federated learning
@@ A.1 Discussion on the Parameter Space
where the $i$-th client will be shrunk to the global with strength $\lambda_{i}$. Therefore, solving (19) will consider different personalization degree for different clients. Establish the minimax statistical error bound for Problem (19) would be open for future work.
Under Assumptions 1 and 2 as we assume, statistical heterogeneity in parameter space (5) would be equivalent to many other similar assumptions for statistical heterogeneity, like $B$-dissimilarity (Li et al., 2020), parameter difference (Chen et al., 2023c) and gradient diversity (T Dinh et al., 2020; Deng et al., 2020).
A.2 Useful Lemmas
To facilitate our analysis, we first present two important lemmas and provide their proof.
@@ A.2 Useful Lemmas
where in the last step, we used the Cauchy inequality.
In addition, notice that ${\tilde{\bm{w}}^{(i)}}$ is the optimal solution for the local model ${{\bm{w}}^{(i)}}$ in Problem (2), thus we have
@@
as claimed. ∎
Remark: Lemma 3 is crucial to bridge the solution of the global model with the local model in Problem 2. Note that the difference is bounded by the gradient norm of local loss functions evaluated at the optimal solution of the global model, which would be decomposed further and controlled by the global model’s statistical accuracy, as we will show later.
@@
Under Assumption 1, 2, 3, for $\tilde{\bm{w}}_{\text{GT}}$, the solution of the GlobalTrain problem in (4), we have
@@
$\leq\frac{\Big{\|}\sum_{i\in[m]}p_{i}\nabla L_{i}({{\bm{w}}_{\star}^{(i)}},S_{i})\Big{\|}+\sum_{i\in[m]}p_{i}\Big{\|}\nabla L_{i}({{\bm{w}}_{\star}^{(g)}},S_{i})-\nabla L_{i}({{\bm{w}}_{\star}^{(i)}},S_{i})\Big{\|}}{\mu/2}.$ (22)
@@
as claimed. ∎
Remark: Lemma 4 implies that if we solve the GlobalTrain problem exactly, the solution will be close to the ground truth of the global model with a certain error, which is independent of $\lambda$. After presenting the two useful lemmas in Appendix A.2, we will start to prove the theoretical results in Section 4.1. In Section A.3, we will prove Theorem 1 first, then we will prove the statistical error bound in Problem 2 as demonstrated in Corollary 1.
A.3 Proof of Statistical Convergence in Theorem 1
In the discuss below, we always assume ${{\bm{w}}_{\star}^{(i)}}$ discussed below comes from the parameter space (5). As the local model is shrunk towards the global model under the influence of personalization degree, how fast the global model ${\tilde{\bm{w}}^{(g)}}$ convergence to ${{\bm{w}}_{\star}^{(g)}}$ can also influence the convergence rate of the local model. In Appendix A.3.1, we will establish the statistical error bound of the global model first. Then in Appendix A.3.2, we will provide the local model error bound, where we explicitly show that the statistical accuracy of the local model depends on the statistical accuracy of the global model. Finally, in Appendix A.3.3, we combine the results of local models and the global model to establish the one-line rate for the local model explicitly stated in Theorem 1.
A.3.1 Global Statistical Error Bound
To analyze the statistical error bound, we consider two separate cases: $R>\sqrt{\sum_{i}\frac{p_{i}}{n_{i}}}$ and $R\leq\sqrt{\sum_{i}\frac{p_{i}}{n_{i}}}$. The motivation for this distinction lies in whether collaboration among different clients is beneficial.
@@ A.3.1 Global Statistical Error Bound
Recall that ${\tilde{\bm{w}}^{(g)}}$ and $\{{\tilde{\bm{w}}^{(i)}}\}_{i\in[m]}$ are the minimizers of Problem (2). According to the first-order condition, we have
@@
For the first term on the R.H.S. in (26), apply the $\mu$-strongly convexity of the loss function $\ell$ (cf. Assumption 2), we then have
@@
Recall the definition in (18) and the assumption about statistical heterogeneity in (5)
@@
If we assume $p_{i}=\frac{1}{m}$ and $n_{i}=n_{j}\forall i\neq j$ (assumptions used in Theorem 1), we have
$\mathbb{E}\left\|\tilde{\Delta}^{(g)}_{\text{stat}}\right\|^{2}\leq\frac{1}{u^{2}}\left(\frac{4\rho^{2}}{n}+2\mu\lambda R^{2}\right).$ (35)
@@
$\sum_{i}p_{i}\nabla L_{i}(\tilde{\bm{w}}_{\text{GT}},S_{i})=0.$ (42)
@@
Take expectation w.r.t all data and use Assumption 3 in the parameter space (5), then we have
@@
If we assume $p_{i}=\frac{1}{m}$, $n_{i}=n_{j}\;\forall i\neq j$ (assumptions used in Theorem 1) and $g(\lambda)\leq 1/(8L)^{2}$, we have
$\mathbb{E}[\|\tilde{\Delta}^{(g)}_{\text{stat}}\|^{2}]$ $\leq\frac{g(\lambda)}{1-4L^{2}g(\lambda)}\left(4L^{2}R^{2}+2\frac{\rho^{2}}{n}\right)+\frac{16\rho^{2}}{\mu^{2}}\frac{1}{N}+\frac{16L^{2}}{\mu^{2}}R^{2}.$ (52)
A.3.2 Local Statistical Error Bound
Analogous to the analysis of the global model, we examine the statistical accuracy of the local model in a similar manner. For simplicity, in the following arguments, we denote the upper bound of $\mathbb{E}\|\tilde{\Delta}^{(g)}_{\text{stat}}\|_{2}$ as $U_{0}$. The expliciate expression would be found in results (35) for $R>\sqrt{\sum_{i}\frac{p_{i}}{n_{i}}}$ and (52) for $R\leq\sqrt{\sum_{i}\frac{p_{i}}{n_{i}}}$.
@@ A.3.2 Local Statistical Error Bound
For the first term on the R.H.S., applying the $\mu$-strongly convex of the loss function $\ell$ (cf. Assumption 2) yields
@@
Plugging (55) back into (54), it yields
@@
Once we denote the upper bound of $\mathbb{E}\left\|\tilde{\Delta}^{(g)}_{\text{stat}}\right\|^{2}$ as $U_{0}$, it yields
@@
If we assume $p_{i}=\frac{1}{m}$, $n_{i}=n_{j}\;\forall i\neq j$ (assumptions used in Theorem 1), we can obtain
$\mathbb{E}\left\|\tilde{\Delta}^{(i)}_{\text{stat}}\right\|^{2}\leq\frac{1}{u^{2}}\left(\frac{4\rho^{2}}{n}+4u\lambda\left(U_{0}^{2}+R^{2}\right)\right).$ (64)
@@
If we assume $p_{i}=\frac{1}{m}$, $n_{i}=n_{j}\;\forall i\neq j$ (assumptions used in Theorem 1), we can obtain
@@
$\overset{(b)}{\leq}\frac{12}{(\mu+\lambda)^{2}}\left[4L^{2}\mathbb{E}[\|{\tilde{\bm{w}}^{(g)}}-{{\bm{w}}_{\star}^{(g)}}\|^{2}]+4L^{2}R^{2}+2\frac{\rho^{2}}{n}\right]+3\mathbb{E}[\|{\tilde{\bm{w}}^{(g)}}-{{\bm{w}}_{\star}^{(g)}}\|^{2}]+3R^{2},$ (65)
where step (a) comes from the assumption of parameter space in (5) and Lemma 3 and step (b) comes from result (A.3.1).
A.3.3 Proof the Statistical Error Bound in Theorem 1
For $R>\sqrt{\sum_{i}\frac{p_{i}}{n_{i}}}$, putting the statistical error of global model (35) into the statistical error of local models (64), we have
@@ A.3.3 Proof the Statistical Error Bound in Theorem 1
$\mathbb{E}\left\|\tilde{\Delta}^{(i)}_{\text{stat}}\right\|^{2}\leq\left(\frac{4\rho^{2}}{u^{2}}+\frac{4\rho^{2}}{u^{3}}\lambda\right)\frac{1}{n}+\frac{8}{u^{2}}\lambda^{2}R^{2}+\frac{4}{u}\lambda R^{2},$ (66)
therefore the second part of (8) in Theorem 1 is established. For clarity, we can define
$q_{2}(\lambda)=\frac{8}{u^{2}}\lambda^{2}R^{2}+\frac{4}{u}\lambda R^{2}.$ (67)
For $R>\sqrt{\sum_{i}\frac{p_{i}}{n_{i}}}$, plugging the result (52) of the global model into the local model’s error bound (A.3.2) and reorganizing the terms yields
$\mathbb{E}\left[\|\tilde{\Delta}^{(i)}_{\text{stat}}\|^{2}\right]\leq$ $\frac{12}{(\mu+\lambda)^{2}}\left[4L^{2}R^{2}+2\frac{\rho^{2}}{n}\right]+{\left[\frac{48L^{2}}{(\mu+\lambda)^{2}}+3\right]U_{0}^{2}+3R^{2}}$ (68)
$\leq$ $\frac{48\rho^{2}}{\mu^{2}}\frac{1}{N}+\left[\frac{48L^{2}}{\mu^{2}}+3\right]R^{2}+\frac{1}{q_{1}(\lambda)},$ (69)
where
$\begin{split}[q_{1}(\lambda)]^{-1}=&\frac{12}{(\mu+\lambda)^{2}}\left[4L^{2}R^{2}+2\frac{\rho^{2}}{n}\right]+\frac{48L^{2}}{(\mu+\lambda)^{2}}\left[\frac{g(\lambda)}{1-4L^{2}g(\lambda)}\left(4L^{2}R^{2}+2\frac{\rho^{2}}{n^{2}}\right)\right.\\ &\left.+2\left(\frac{8\rho^{2}}{\mu^{2}}\frac{1}{N}+\frac{8L^{2}}{\mu^{2}}R^{2}\right)\right]+\frac{3g(\lambda)}{1-4L^{2}g(\lambda)}\left(4L^{2}R^{2}+2\frac{\rho^{2}}{n^{2}}\right).\end{split}$ (70)
This combining with the definition of $g(\lambda)$ implies that $q_{1}(\lambda)$ is a monotonically increasing function of $\lambda$ and $\lim_{\lambda\to\infty}q_{1}(\lambda)=\infty$, as claimed.
A.4 The Proof of Corollary 1
Analogous to the proof of Theorem 1, we will discuss the two cases of $R^{2}$ separately as well. For clarity of our analysis, we will assume $p_{i}=\frac{1}{m}$ and $n_{i}=n_{j}\>\forall i\neq j$.
A.4.1 The Statistical Convergence of the Global Model with a Choice of $\lambda$
Case 1: For $R>\frac{1}{\sqrt{n}}$, we have derived the global model error bound in (34) as
If we set the $\lambda$ as
$\lambda\leq\frac{\rho^{2}}{\mu nR^{2}},$ (72)
then solving the inequality in (35) w.r.t $\sum_{i}p_{i}\mathbb{E}\left\|\tilde{\Delta}^{(i)}_{\text{stat}}\right\|^{2}$ cound yield
$\begin{split}\mathbb{E}\left\|\tilde{\Delta}^{(g)}_{\text{stat}}\right\|\leq\left(\sum_{i}p_{i}\mathbb{E}\left\|\tilde{\Delta}^{(i)}_{\text{stat}}\right\|^{2}\right)^{\frac{1}{2}}\leq\frac{(1+\sqrt{3})\rho}{\mu\sqrt{n}}.\end{split}$ (73)
Case 2: For $R\leq\frac{1}{\sqrt{n}}$, we have obtain the results in (52) as
$\mathbb{E}[\|\tilde{\Delta}^{(g)}_{\text{stat}}\|^{2}]$ $\leq\frac{g(\lambda)}{1-4L^{2}g(\lambda)}\left(4L^{2}R^{2}+2\frac{\rho^{2}}{n}\right)+\frac{16\rho^{2}}{\mu^{2}}\frac{1}{N}+\frac{16L^{2}}{\mu^{2}}R^{2}.$ (74)
If we assume
we have
Furthermore, we can get a simplified condition for $\lambda$ as
$\lambda\geq\frac{8\kappa}{\mu}\left\{8L^{2}\vee\left(\left(2L^{2}R^{2}+\frac{\rho^{2}}{n}\right)\Big{/}\left(\frac{4\rho^{2}}{\mu^{2}N}+\frac{4L^{2}R^{2}}{\mu^{2}}\right)\right)\right\}-\mu,$ (77)
Combine two-case arguments, we have
$\mathbb{E}\|{\tilde{\bm{w}}^{(g)}}-{{\bm{w}}_{\star}^{(g)}}\|^{2}\leq\begin{cases}\left(\frac{(1+\sqrt{3})\rho}{\mu\sqrt{n}}\right)^{2}&R>\frac{1}{\sqrt{n}}\\ 32\left(\frac{\rho^{2}}{\mu^{2}}\frac{1}{N}+\frac{L^{2}}{\mu^{2}}R^{2}\right)&R\leq\frac{1}{\sqrt{n}}.\end{cases}$ (78)
Thus it yields a one-line rate
$\mathbb{E}\|{\tilde{\bm{w}}^{(g)}}-{{\bm{w}}_{\star}^{(g)}}\|^{2}\leq C_{1}\frac{1}{N}+C_{2}(\frac{1}{n}\wedge R^{2}),$ (79)
where
A.4.2 The Statistical Convergence of the Local Model with a Choice of $\lambda$
Case 1: For $R>\frac{1}{\sqrt{n}}$, we have derived the result in (64) as
$\mathbb{E}\left\|\tilde{\Delta}^{(i)}_{\text{stat}}\right\|^{2}\leq\frac{1}{u^{2}}\left(\frac{4\rho^{2}}{n}+4u\lambda\left(U_{0}^{2}+R^{2}\right)\right).$ (81)
In addition, recall that, when $R>\frac{1}{\sqrt{n}}$, we set $\lambda\leq\frac{\rho^{2}}{\mu nR^{2}}$ in the statistical convergence analysis of the global model. Therefore, putting the global model’s error bound (See result in (78)) into (64), which yields
$\begin{split}\mathbb{E}\|\tilde{\Delta}^{(i)}_{\text{stat}}\|_{2}&\leq\frac{1}{\mu}\left[\frac{\rho}{\sqrt{n}}+\left(\frac{\rho^{2}}{n}+\frac{2\rho^{2}}{nR^{2}}\left(U_{0}+R\right)^{2}\right)^{\frac{1}{2}}\right].\\ &=\frac{1}{\sqrt{n}}\frac{1}{\mu}\left(\rho+\left(\rho^{2}+\frac{2\rho^{2}}{R^{2}}(U_{0}+R)^{2}\right)^{\frac{1}{2}}\right)\\ &\leq\frac{1}{\sqrt{n}}\frac{\rho+(\rho^{2}+4\rho^{2}(C_{2}+1))^{\frac{1}{2}}}{\mu}\\ &=\frac{1}{\sqrt{n}}\frac{\rho(1+(1+4(C_{2}+1))^{\frac{1}{2}})}{\mu},\end{split}$ (82)
where $C_{2}$ has been defined in (80).
Case 2: For $R\leq\frac{1}{\sqrt{n}}$, from the result (A.3.2), we know
$\quad\mathbb{E}[\|{\tilde{\bm{w}}^{(i)}}-{{\bm{w}}_{\star}^{(i)}}\|^{2}]\leq\frac{12}{(\mu+\lambda)^{2}}\left[4L^{2}U_{0}^{2}+4L^{2}R^{2}+2\frac{\rho^{2}}{n}\right]+3U_{0}^{2}+3R^{2}.$ (83)
Then putting the global model’s results (78) into (A.3.2), it yields
$\quad\mathbb{E}[\|{\tilde{\bm{w}}^{(i)}}-{{\bm{w}}_{\star}^{(i)}}\|^{2}]$ (84)
$\leq\frac{12}{(\mu+\lambda)^{2}}\left[128L^{2}\left(\frac{\rho^{2}}{\mu^{2}}\frac{1}{N}+\frac{L^{2}}{\mu^{2}}R^{2}\right)+4L^{2}R^{2}+2\frac{\rho^{2}}{n}\right]+96\left(\frac{\rho^{2}}{\mu^{2}}\frac{1}{N}+\frac{L^{2}}{\mu^{2}}R^{2}\right)+3R^{2}$
$=\frac{12}{(\mu+\lambda)^{2}}\left[\frac{128L^{2}\rho^{2}}{\mu^{2}}\frac{1}{N}+(128L^{4}/\mu^{2}+4L^{2})R^{2}+2\frac{\rho^{2}}{n}\right]+\frac{96\rho^{2}}{\mu^{2}}\frac{1}{N}+(96\frac{L^{2}}{\mu^{2}}+3)R^{2}.$ (85)